i cnt’t solve this trigo. problem.....
please give me hints to solve it....
A+B+C=180
tanA+ tanB+ tanC=0
cotA+cotB+cotC=?

Grade:11

1 Answers

ABHISHAY TURNER
6 years ago
Given- A+B+C=180=>A+B=180 – CMULTIPLYING COT ON BOTH SIDES=>COT(A+B)=COT(180-C) =>COT(A+B)= ––––––( 1 ) =>COT(A+B)= 1 – COT C => – COT C = 1 – COT C HENCE,COT A+ COT B+ COT C=1